OpenStudy (thadds2003):

sorry if i sound dumb, because i suck at fractions, what is 4/5 0f 30?

OpenStudy (alexandervonhumboldt2):

4/5*30=(30*4)/5=? or 4/5*30=4*6=?

OpenStudy (thadds2003):

whaaaaaa?

jabez177 (jabez177):

\[4\div5\times30=?\]

OpenStudy (thadds2003):

24?

jabez177 (jabez177):

Yes. :D

OpenStudy (sam141101):

five divided by 30 =6 so 6+6+6+6=?

OpenStudy (thadds2003):

so thats the answer?

OpenStudy (alexandervonhumboldt2):

yes it is 24

OpenStudy (rainbowdashrocks):

@Thadds2003

OpenStudy (thadds2003):

huh?

OpenStudy (rainbowdashrocks):

you didnt sound dumb ya know

OpenStudy (thadds2003):

i thought i did, its a lame question to ask. XD!

OpenStudy (rainbowdashrocks):

i dont even know so its not lame and sorry i couldnt help busy

jabez177 (jabez177):

Don't worry, we're all learning! :D

OpenStudy (rainbowdashrocks):

yup